# Quantum synchronization between two strongly driven YIG spheres mediated via a microwave cavity

### v2

@@ III Results
We first choose $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}\Omega_{1}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\bar{q}_{1}$, $\bar{q}_{2}$ and $\bar{p}_{1}$, $\bar{p}_{2}$ become indistinguishable at times as long as $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of two YIG spheres have been previously explored in a cavity-magnomechanical setup, in which the phase synchronization has been achieved between the vibrational degrees of freedom, instead of magnon modes PhysRevResearch.5.043197. When the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, the system still maintains the phase synchronization, however with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f). Note that at the steady state, though the quadrature variables vary with time, the phase $\phi$ remains constant.
We further consider that the coupling strengths $g_{i}$ of these magnon modes with the cavity are different from each other. This could arise when the YIG spheres are trapped inside the cavity in an asymmetrical configuration with respect to the central antinodes of the cavity. When we choose $g_{1}>g_{2}$, the amplitude of the second magnon mode becomes lower compared to that of the first. Nevertheless, the system maintains phase synchronization with a phase difference of $\phi=0.1320$ radian (Fig. 2g, 2h and 2i). We note that such a phase relationship between coupled magnonic oscillators can be continuously tuned by varying coupling strengths PhysRevResearch.5.043197; PhysRevResearch.6.033207.
@@ IV Conclusion
Second, and more intriguingly, we observed the emergence of quantum synchronization between the coupled magnon modes (Fig. 4). This phenomenon is quantitatively characterized by the synchronization measure $S_{q}^{\phi}$, which approaches unity under optimal conditions. In this regime, the individual quantum fluctuations of each mode become phase-locked, leading to coherent evolution despite the presence of decoherence and thermal noise. This indicates that the system dynamically suppresses relative quantum uncertainties, allowing for robust synchronization at the quantum level. In addition, we explored the effect of temperature on this delicate quantum behavior (Fig. 4). We find that as thermal noise increases, the degree of quantum synchronization gradually decreases. This finding underscores the importance of low-temperature environments when trying to observe and maintain quantum synchronization in experiments.
@@ V ACKNOWLEDGMENTS
$P_{i}$ $=-\frac{\gamma_{i}}{2}+\iota\Bigg[A_{i}g_{i}\alpha_{i}\beta^{*}\!\left(2-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-C_{i})}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+C_{i})}$
$\hskip 17.00024pt-g_{i}A_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt+A_{i}\Omega_{i}\alpha_{i}e^{-\iota t(D_{i}+C_{i})}\left(2-A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)\Bigg]\;,$
$Q_{i}$ $=\iota\Bigg[-A_{i}^{2}g_{i}\alpha_{i}^{3}\beta^{*}-A_{i}\Omega_{i}\alpha_{i}e^{\iota t(D_{i}+C_{i})}-A_{i}g_{i}\alpha_{i}\beta e^{-\iota t(B_{i}-C_{i})}$
$\hskip 17.00024pt-A_{i}\Omega_{i}\alpha_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(D_{i}+C_{i})}-A_{i}^{2}\Omega_{i}\alpha_{i}^{3}e^{-\iota t(D_{i}+C_{i})}$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}\beta\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\Bigg]\;,$
$R_{i}$ $=-\iota g_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\;,$
$S_{i}$ $=\iota g_{i}A_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;,$
$U_{i}$ $=-\iota g_{i}\left(1-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-C_{i})}\;,$
$W_{i}$ $=\iota A_{i}g_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;,$

### v3

@@ III Results
We first choose $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}\Omega_{1}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\bar{q}_{1}$, $\bar{q}_{2}$ and $\bar{p}_{1}$, $\bar{p}_{2}$ become indistinguishable at times as long as $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of two YIG spheres have been previously explored in a cavity-magnomechanical setup, in which the phase synchronization has been achieved between the vibrational degrees of freedom, instead of magnon modes PhysRevResearch.5.043197 . When the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, the system still maintains the phase synchronization, with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f) between the trajectories of the two magnon modes, although the quadratures still oscillate at the same frequency. Such a behavior refers to a classical phase synchronization.
We further consider that the coupling strengths $g_{i}$ of these magnon modes with the cavity are different from each other. This could arise when the YIG spheres are trapped inside the cavity in an asymmetrical configuration with respect to the central antinodes of the cavity. When we choose $g_{1}>g_{2}$, the amplitude of the second magnon mode becomes lower compared to that of the first. Nevertheless, the system maintains phase synchronization with a phase difference of $\phi=1.3357$ radian (Fig. 2g, 2h and 2i). We note that such a phase relationship between coupled magnonic oscillators can be continuously tuned by varying coupling strengths PhysRevResearch.5.043197 ; PhysRevResearch.6.033207 .
@@
In contrast Figure 4a, for a small thermal population $\bar{n}_{m}=0.1$, $S_{q}^{\phi}$ still grows monotonically but saturates at a reduced value $S_{q}^{\phi}\simeq 0.83$, reflecting partial degradation of synchronization due to thermal noise. Nevertheless, in both cases the system reaches a steady synchronized regime, demonstrating that cavity-mediated magnon synchronization remains robust against weak thermal decoherence.
@@ IV Conclusion
Second, and more intriguingly, we observed the emergence of quantum synchronization between the coupled magnon modes (Fig. 4b). This phenomenon is quantitatively characterized by the synchronization measure $S_{q}^{\phi}$, which approaches unity under optimal conditions. In this regime, the individual quantum fluctuations of each mode become phase-locked, leading to coherent evolution despite the presence of decoherence and thermal noise. This indicates that the system dynamically suppresses relative quantum uncertainties, allowing for robust synchronization at the quantum level. In addition, we explored the effect of temperature on this delicate quantum behavior (Fig. 4b). We find that as thermal noise increases, the degree of quantum synchronization gradually decreases. This finding underscores the importance of low-temperature environments when trying to observe and maintain quantum synchronization in experiments.
@@ V ACKNOWLEDGMENTS
$P_{i}$ $=-\frac{\gamma_{i}}{2}+\iota\Bigg[A_{i}g_{i}\alpha_{i}\beta^{*}\!\left(2-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-\langle C_{i}\rangle)}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+\langle C_{i}\rangle)}$
$\hskip 17.00024pt-g_{i}A_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-\langle C_{i}\rangle)}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+\langle C_{i}\rangle)}\left(1+A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt+A_{i}\Omega_{i}\alpha_{i}e^{-\iota t(D_{i}+\langle C_{i}\rangle)}\left(2-A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-\langle C_{i}\rangle)}\left(1+A_{i}|\alpha_{i}|^{2}\right)\Bigg]\;,$
$Q_{i}$ $=\iota\Bigg[-A_{i}^{2}g_{i}\alpha_{i}^{3}\beta^{*}-A_{i}\Omega_{i}\alpha_{i}e^{\iota t(D_{i}+\langle C_{i}\rangle)}-A_{i}g_{i}\alpha_{i}\beta e^{-\iota t(B_{i}-\langle C_{i}\rangle)}$
$\hskip 17.00024pt-A_{i}\Omega_{i}\alpha_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(D_{i}+\langle C_{i}\rangle)}-A_{i}^{2}\Omega_{i}\alpha_{i}^{3}e^{-\iota t(D_{i}+\langle C_{i}\rangle)}$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}\beta\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-\langle C_{i}\rangle)}\Bigg]\;,$
$R_{i}$ $=-\iota g_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-\langle C_{i}\rangle)}\;,$
$S_{i}$ $=\iota g_{i}A_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-\langle C_{i}\rangle)}\;,$
$U_{i}$ $=-\iota g_{i}\left(1-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-\langle C_{i}\rangle)}\;,$
$W_{i}$ $=\iota A_{i}g_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-\langle C_{i}\rangle)}\;,$